# Image Segmentation with Large Language Models: A Survey with Perspectives for Intelligent Transportation Systems

### v1

@@ Abstract
publicationid: pubid: $0$–$00$ 2025 IEEE
@@ I Introduction
In ITS, LLM-augmented VLSeg has transformative potential across several applications. For autonomous driving, it enables real-time scene parsing, dynamic obstacle detection, and predictive segmentation of crash scenarios, as demonstrated by frameworks like DriveLM [23] and InsightGPT [24]. For traffic monitoring, VLSeg supports smarter traffic flow analysis and anomaly detection in surveillance feeds, enhancing city-scale mobility solutions [25, 26]. Datasets like Talk2Car [27] and Road-Seg-VL [28] provide language-guided annotations, facilitating the development of models that align visual perception with human instructions [1]. Moreover, recent advancements in open-vocabulary segmentation, such as CLIPSeg [29] and OpenSeg [30], enable zero-shot segmentation of unseen objects, addressing the open-world challenges inherent in ITS [30].
@@ II-B1 Historical Evolution of Segmentation Approaches
The first major breakthrough in deep learning-based segmentation was the Fully Convolutional Network (FCN) [43], which replaced the dense, fully-connected layers of classification networks (like AlexNet [53]) with convolutional layers, enabling end-to-end training for pixel-wise prediction. For ITS, this meant that models could now learn to segment entire driving scenes at once. Following this, architectures like SegNet [44] and U-Net [45] introduced the powerful encoder-decoder paradigm. The encoder, typically a pre-trained classification network (e.g., VGG [54]), progressively downsamples the input to capture semantic context, while the decoder upsamples these features to reconstruct a full-resolution segmentation map. U-Net’s key innovation was the use of ”skip connections,” which pass fine-grained feature details from the encoder directly to the decoder, proving crucial for accurately localizing small objects like traffic signs and pedestrians in ITS scenes. To handle the vast variation in object scales in driving environments (e.g., distant cars vs. nearby trucks), the DeepLab family of models [3, 9] introduced atrous (or dilated) convolutions and Atrous Spatial Pyramid Pooling (ASPP), allowing the network to probe features at multiple resolutions without increasing computational cost. These CNN-based models became the dominant approach and achieved impressive results on benchmarks like Cityscapes [7].
@@ III Architectural Deep Dive
Input ImageText PromptVision EncoderLanguage EncoderMulti-ModalFusion ModuleMask DecoderSegmentationMask
@@ III-A The Vision Encoder Backbone
  • Vision Transformers (ViTs): The introduction of the Vision Transformer [12] marked a paradigm shift. By treating an image as a sequence of patches and applying self-attention, ViTs can model long-range dependencies across the entire image. This is highly advantageous for VLSeg, as a prompt like ”segment the car farthest away” requires a global understanding of the scene. The original ViT architecture produces a single-resolution feature map, which can be suboptimal for dense prediction tasks like segmentation.
@@ III-B The Language Encoder and Prompt Engineering
  • General-Purpose LLMs: Other approaches integrate more general-purpose language models, such as BERT [15] or T5 [17]. While not specifically pre-trained for vision-language alignment, these models often have a deeper understanding of syntax and relational language, which can be beneficial for interpreting complex, compositional prompts (e.g., ”segment the second car to the left of the traffic light”).
  • Prompt Engineering: The performance of a VLSeg model can be surprisingly sensitive to the phrasing of the text prompt. This has given rise to the sub-field of ”prompt engineering.” Research has shown that providing more descriptive prompts often yields better results. For example, instead of ”car”, using ”the red sports car” helps the model better disambiguate objects. Furthermore, some models benefit from ”prompt tuning” or ”prompt learning,” where a small set of learnable embedding vectors are prepended to the text prompt. These vectors are optimized during training to steer the language encoder towards producing embeddings that are more effective for the downstream segmentation task, without needing to fine-tune the entire large language model [68, 69].
@@ III-C The Mask Decoder
SAMMobileSAMSEEMCLIPSegOpenSeg005005005005001,00010001{,}0001 ,  [^0] Value Inference Time (ms)Model Size (MB)
@@ IV-A Categorization by Prompting Interface
  • Multi-modal and Interactive Prompts: The most advanced models offer a combination of prompt types for maximum flexibility. SEEM (Segment Everything Everywhere Model) [21] is a prime example, accepting text, points, boxes, scribbles, and even other image masks as prompts. This allows for a rich, interactive dialogue between the user and the system, enabling complex instructions and iterative refinement. In ITS, this could translate to a system that can take an initial command (”segment all vehicles”) and then refine it with a point prompt (”…but only this one”).
@@ IV-B Categorization by Core Architecture
  • Promptable Foundation Models: This category is defined by SAM [31]. The architecture consists of a powerful image encoder (ViT-H), a flexible prompt encoder, and a lightweight mask decoder. Its key innovation is being pre-trained for the task of ”promptable segmentation” on a massive dataset (SA-1B). While SAM itself has limited text understanding, its architecture is designed to be extended.
@@ IV-C Multi-Modal Fusion Strategies
  • Advanced and Hybrid Fusion: To find a better balance, more advanced techniques have been developed. Gated fusion introduces mechanisms that learn to control the flow of information from each modality, deciding how much visual or textual information to use at different processing stages. Bilinear pooling offers a much richer fusion by capturing every pairwise interaction between the visual and language features, but this is often too computationally expensive for real-time applications [72, 73]. Consequently, many state-of-the-art models use hybrid approaches. They might employ self-attention within each modality to create refined feature representations first, followed by multiple layers of cross-attention to achieve a deep, iterative alignment between vision and language before the final decoding step [74].
@@ V A Comparative Analysis of State-of-the-Art Models
Grounded-SAM [20] SAM + Grounding DINO Text, Box Combines open-vocabulary detection with SAM Segmenting objects based on descriptive language (e.g., ”the red car turning left”).
@@
A key architectural trade-off is illustrated by comparing hybrid models like Grounded-SAM with unified models like SEEM. Grounded-SAM, by chaining a specialized open-vocabulary detector (Grounding DINO) with a powerful segmenter (SAM), excels at tasks where the primary challenge is to first *find* a specific object in a cluttered scene based on a descriptive prompt. Its failure modes often stem from the detector failing; if the object isn’t found, it can’t be segmented. In contrast, SEEM’s unified architecture is more flexible, handling a wider array of prompts (scribbles, points) and potentially performing better at segmenting amorphous regions (e.g., ”segment the puddle”) that lack clear object boundaries for a detector. Its weakness may lie in handling highly complex compositional prompts, where the chained reasoning of a detector-segmenter might be more robust. The choice between these approaches depends on the specific ITS task: for finding and segmenting known categories of objects with high precision, the hybrid approach is strong; for interactive annotation and flexible human-AI collaboration, the unified model offers advantages.
@@ V-A Quantitative Performance Benchmarking
While qualitative comparisons are useful for understanding architectural innovations, quantitative benchmarks are essential for evaluating practical performance. Table III presents a summary of reported performance metrics for several key VLSeg models on standard ITS-relevant datasets. It is important to note that direct comparisons can be challenging due to variations in experimental setups, such as the specific vocabulary used for open-set evaluation or whether the model was fine-tuned on the target dataset.
@@
The results in Table III reveal several key trends. First, models specifically designed and optimized for a single task and dataset, like OneFormer on Cityscapes panoptic segmentation, still often outperform more general, open-vocabulary models in their specific domain. This highlights a critical trade-off that can be termed the ”cost of generalization.” The high Panoptic Quality (PQ) of OneFormer (68.0) compared to the scores of models designed for open-ended tasks demonstrates that there is currently a performance penalty for the flexibility that VLSeg provides. This is particularly true for safety-critical sub-tasks like robust sidewalk detection, where specialized ensemble models have been shown to surpass the capabilities of more general LLM-based approaches [77]. While models like OpenSeg show promising open-vocabulary mIoU on complex urban scenes, they do not yet match the performance of closed-set, specialized systems. This gap suggests that for safety-critical ITS applications requiring the highest possible accuracy on a known set of classes (e.g., standard traffic signs, lane markings), specialized models remain superior. However, for handling novelty and improving human-AI interaction, the ”cost” of using a more general VLSeg model is justifiable.
@@ V-A2 Standardizing Evaluation for VLSeg in ITS
    ITS-Specific Vocabulary: Evaluation should use a standardized set of ITS-specific prompts, including both common categories (e.g., ”car,” ”pedestrian”) and more complex, compositional queries (e.g., ”red car turning left,” ”pedestrian crossing between parked vehicles”).
@@
    Robustness Evaluation: Performance under corruptions (e.g., motion blur, adverse weather) should be systematically evaluated using standardized corruption sets like Cityscapes-C or BDD100K-C.
@@ VI Advanced and Emerging Topics
3D VLSeg presents unique challenges. Point clouds are sparse, unordered, and unstructured, making them fundamentally different from the dense, grid-like structure of images. Early work in 3D segmentation focused on adapting CNN-like architectures to operate on voxels or directly on points [80, 81]. More recently, the focus has shifted to grounding language in these 3D spaces. Models like LidarCLIP [82] learn to align text descriptions with entire 3D point clouds. Building on this, 3D VLSeg models aim to segment specific objects in the point cloud based on a text prompt. This is often achieved by projecting 2D image features (from multiple camera views) onto the 3D point cloud, creating a text-aware 3D representation. The system can then respond to queries like ”segment the point cloud of the truck in front of us” or ”highlight the curb on the right.” This allows for a much more intuitive and powerful way to interact with and understand 3D sensor data, which is essential for tasks like 3D object detection, motion prediction, and path planning [83, 84].
@@ VI-B Video and Temporal Consistency
Several approaches are being explored. One common method is to use optical flow to propagate masks from one frame to the next. However, this can be error-prone, especially with fast-moving objects or camera motion. A more robust approach, exemplified by models like XMem [61], is to use a memory-based architecture. In this paradigm, the model maintains a ”memory” of past frames and their segmentation masks. When processing a new frame, it uses attention mechanisms to query this memory, allowing it to re-identify and maintain a consistent segmentation of objects over long video sequences. Language can be used to initialize the tracking (e.g., ”start tracking the blue car”) and to re-identify objects if the tracking is lost (”where is the blue car now?”). As video foundation models become more powerful, we expect to see more end-to-end V-VLSeg models that can reason about actions and events over time [85, 86].
@@ VI-C Federated and Collaborative Learning
Training powerful segmentation models requires vast amounts of diverse data, which raises significant privacy concerns, especially when the data is collected from personal vehicles. Federated Learning (FL) is a machine learning paradigm that addresses this issue. Instead of pooling raw data in a central server, the central model is sent to individual vehicles (or ”clients”). Each client updates the model locally using its own private data, and only the model updates (gradients or weights) are sent back to the server to be aggregated. This allows a global model to learn from the collective data of the entire fleet without any raw driving data ever leaving the vehicle [87, 88]. Applying FL to VLSeg in ITS is an active research area, focusing on challenges like communication efficiency and handling the non-IID (non-independently and identically distributed) nature of data from different vehicles [89].
A related concept is Collaborative (or Collective) Perception. In this scenario, vehicles and infrastructure (e.g., smart traffic lights) communicate with each other, sharing high-level perception information, such as segmentation masks or object detections. For example, a vehicle whose view is occluded by a large truck could receive segmentation data from another vehicle at a better vantage point, allowing it to ”see” the pedestrian that is about to cross the street. This creates a more robust and complete understanding of the driving scene than any single agent could achieve alone. Research in this area focuses on what information to share, how to fuse it effectively, and how to ensure the communication is secure and reliable [90, 91]. Language can act as a powerful and efficient communication medium in these systems, where one agent could send a compressed, semantic message like ”pedestrian crossing from your right” to another.

### v2

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
@@ I Introduction
In ITS, LLM-augmented VLSeg has transformative potential across several applications. For autonomous driving, it enables real-time scene parsing, dynamic obstacle detection, and predictive segmentation of crash scenarios, as demonstrated by frameworks like DriveLM [23] and InsightGPT [24]. For traffic monitoring, VLSeg supports smarter traffic flow analysis and anomaly detection in surveillance feeds, enhancing city-scale mobility solutions [25, 26]. Datasets like Talk2Car [27] and Road-Seg-VL [28] provide language-guided annotations, facilitating the development of models that align visual perception with human instructions [1]. Moreover, recent advancements in open-vocabulary segmentation, such as CLIPSeg [29] and OpenSeg [30], enable zero-shot segmentation of unseen objects, addressing the open-world challenges inherent in ITS [30]. A central theme of this survey is the critical trade-off we term the ’cost of generalization’: while powerful VLSeg models offer unprecedented flexibility for open-world scenarios, they often lag in performance on specific, safety-critical tasks compared to highly optimized, specialized models. Analyzing this trade-off is essential for their practical deployment in ITS.
@@ II-B1 Historical Evolution of Segmentation Approaches
Figure 2: Timeline of key milestones in the evolution of segmentation for ITS, with a color-coded legend at the top indicating categories.
@@
The first major breakthrough in deep learning-based segmentation was the Fully Convolutional Network (FCN) [43], which replaced the dense, fully-connected layers of classification networks (like AlexNet [53]) with convolutional layers, enabling end-to-end training for pixel-wise prediction. For ITS, this meant that models could now learn to segment entire driving scenes at once. Following this, architectures like SegNet [44] and U-Net [45] introduced the powerful encoder-decoder paradigm. The encoder, typically a pre-trained classification network (e.g., VGG [54]), progressively downsamples the input to capture semantic context, while the decoder upsamples these features to reconstruct a full-resolution segmentation map. U-Net’s key innovation was the use of “skip connections,” which pass fine-grained feature details from the encoder directly to the decoder, proving crucial for accurately localizing small objects like traffic signs and pedestrians in ITS scenes. To handle the vast variation in object scales in driving environments (e.g., distant cars vs. nearby trucks), the DeepLab family of models [3, 9] introduced atrous (or dilated) convolutions and Atrous Spatial Pyramid Pooling (ASPP), allowing the network to probe features at multiple resolutions without increasing computational cost. These CNN-based models became the dominant approach and achieved impressive results on benchmarks like Cityscapes [7].
@@ II-D Related Work and Contributions
Low-Level Research (Core Components & Concepts) Mid-Level Research (Task-Oriented VLSeg Models) High-Level Research (Integrated Systems & Applications) Foundational Algorithms (e.g., FCN, U-Net, DeepLab) Vision Backbones (e.g., ViT, Swin Transformer) V-L Pre-training (e.g., CLIP, DINOv2) Zero-Shot & Open-Vocabulary (e.g., CLIPSeg, OpenSeg) Promptable Foundation Models (e.g., SAM) Hybrid & Interactive Models (e.g., Grounded-SAM, SEEM) End-to-End Driving (e.g., DriveLM, LMDrive) Traffic Management (e.g., TrafficGPT, LLM4Tr) Human-AI Collaboration (e.g., Talk2BEV, Voyager)
Figure 3: A hierarchical categorization of the reviewed literature. Research flows from low-level foundational concepts to mid-level task-specific models, and finally to high-level integrated systems for ITS applications.
@@ III-A The Vision Encoder Backbone
  • Vision Transformers (ViTs): The introduction of the Vision Transformer [12] marked a paradigm shift. By treating an image as a sequence of patches and applying self-attention, ViTs can model long-range dependencies across the entire image. This is highly advantageous for VLSeg, as a prompt like “segment the car farthest away” requires a global understanding of the scene. The original ViT architecture produces a single-resolution feature map, which can be suboptimal for dense prediction tasks like segmentation.
@@ III-B The Language Encoder and Prompt Engineering
  • General-Purpose LLMs: Other approaches integrate more general-purpose language models, such as BERT [15] or T5 [17]. While not specifically pre-trained for vision-language alignment, these models often have a deeper understanding of syntax and relational language, which can be beneficial for interpreting complex, compositional prompts (e.g., “segment the second car to the left of the traffic light”).
  • Prompt Engineering: The performance of a VLSeg model can be surprisingly sensitive to the phrasing of the text prompt. This has given rise to the sub-field of “prompt engineering.” Research has shown that providing more descriptive prompts often yields better results. For example, instead of “car”, using ”the red sports car” helps the model better disambiguate objects. Furthermore, some models benefit from “prompt tuning” or “prompt learning,” where a small set of learnable embedding vectors are prepended to the text prompt. These vectors are optimized during training to steer the language encoder towards producing embeddings that are more effective for the downstream segmentation task, without needing to fine-tune the entire large language model [68, 69].
@@ IV-A Categorization by Prompting Interface
  • Multi-modal and Interactive Prompts: The most advanced models offer a combination of prompt types for maximum flexibility. SEEM (Segment Everything Everywhere Model) [21] is a prime example, accepting text, points, boxes, scribbles, and even other image masks as prompts. This allows for a rich, interactive dialogue between the user and the system, enabling complex instructions and iterative refinement. In ITS, this could translate to a system that can take an initial command (“segment all vehicles”) and then refine it with a point prompt (“…but only this one”).
@@ IV-B Categorization by Core Architecture
  • Promptable Foundation Models: This category is defined by SAM [31]. The architecture consists of a powerful image encoder (ViT-H), a flexible prompt encoder, and a lightweight mask decoder. Its key innovation is being pre-trained for the task of “promptable segmentation” on a massive dataset (SA-1B). While SAM itself has limited text understanding, its architecture is designed to be extended.
@@ IV-C Multi-Modal Fusion Strategies
Early FusionVTConcatFused Advantage: Simple and computationally cheap. Limitation: Semantically weak; poor at resolving complex prompts.
ITS Relevance: Suitable only for simple, non-critical tasks. Cross-AttentionVTAttnFusedQueryKey/Value Advantage: State-of-the-art grounding language in vision. Limitation: High computational cost ($\mathcal{O}(MN)$), latency bottleneck. ITS Relevance: Powerful for understanding, but challenges for real-time use. Gated FusionVTGateFusedControl Advantage: Dynamically controls information flow from each modality.
[…]
Figure 6: Comparison of Multi-Modal Fusion Mechanisms in VLSeg for ITS Applications. Each strategy presents a different trade-off between computational cost, semantic richness, and suitability for real-time systems.
@@
  • Advanced and Hybrid Fusion: To find a better balance, more advanced techniques have been developed. Gated fusion introduces mechanisms that learn to control the flow of information from each modality, deciding how much visual or textual information to use at different processing stages. Bilinear pooling offers a much richer fusion by capturing every pairwise interaction between the visual and language features, but this is often too computationally expensive for real-time applications [72, 73]. Consequently, many state-of-the-art models use hybrid approaches. They might employ self-attention within each modality to create refined feature representations first, followed by multiple layers of cross-attention to achieve a deep, iterative alignment between vision and language before the final decoding step [74]. The clear dominance of cross-attention fusion underscores a fundamental shift in the field: the priority is now on achieving precise semantic grounding of language in vision, even at a high computational cost. For real-time ITS, this presents a critical bottleneck and is the primary motivation for the intense research into efficient attention variants and alternative architectures that can bridge this performance gap without sacrificing semantic understanding.
@@ V A Comparative Analysis of State-of-the-Art Models
Grounded-SAM [20] SAM + Grounding DINO Text, Box Combines open-vocabulary detection with SAM Segmenting objects based on descriptive language (e.g., “the red car turning left”).
@@
A key architectural trade-off is illustrated by comparing hybrid models like Grounded-SAM with unified models like SEEM. Grounded-SAM, by chaining a specialized open-vocabulary detector (Grounding DINO) with a powerful segmenter (SAM), excels at tasks where the primary challenge is to first *find* a specific object in a cluttered scene based on a descriptive prompt. Its failure modes often stem from the detector failing; if the object isn’t found, it can’t be segmented. In contrast, SEEM’s unified architecture is more flexible, handling a wider array of prompts (scribbles, points) and potentially performing better at segmenting amorphous regions (e.g., ”segment the puddle”) that lack clear object boundaries for a detector. Its weakness may lie in handling highly complex compositional prompts, where the chained reasoning of a detector-segmenter might be more robust. The choice between these approaches depends on the specific ITS task: for finding and segmenting known categories of objects with high precision, the hybrid approach is strong; for interactive annotation and flexible human-AI collaboration, the unified model offers advantages. This architectural dichotomy has significant implications for verification and validation; the modular nature of a hybrid system like Grounded-SAM may lend itself more readily to established safety standards like ISO 26262, as individual components (detector, segmenter) could potentially be certified independently. In contrast, the end-to-end nature of a unified model like SEEM might require new validation methodologies to ensure safety-critical reliability.
@@ V-A Quantitative Performance Benchmarking
While qualitative comparisons are useful for understanding architectural innovations, quantitative benchmarks are essential for evaluating practical performance. It is important to note that the performance metrics presented in this section are compiled from the original papers for comparative purposes; direct comparisons can be challenging due to variations in experimental setups and evaluation protocols. Table III presents a summary of reported performance metrics for several key VLSeg models on standard ITS-relevant datasets. It is important to note that direct comparisons can be challenging due to variations in experimental setups, such as the specific vocabulary used for open-set evaluation or whether the model was fine-tuned on the target dataset.
@@
The results in Table III reveal several key trends. First, models specifically designed and optimized for a single task and dataset, like OneFormer on Cityscapes panoptic segmentation, still often outperform more general, open-vocabulary models in their specific domain. This highlights a critical trade-off that can be termed the “cost of generalization.” The high Panoptic Quality (PQ) of OneFormer (68.0) compared to the scores of models designed for open-ended tasks demonstrates that there is currently a performance penalty for the flexibility that VLSeg provides. This is particularly true for safety-critical sub-tasks like robust sidewalk detection, where specialized ensemble models have been shown to surpass the capabilities of more general LLM-based approaches [77]. While models like OpenSeg show promising open-vocabulary mIoU on complex urban scenes, they do not yet match the performance of closed-set, specialized systems. This gap suggests that for safety-critical ITS applications requiring the highest possible accuracy on a known set of classes (e.g., standard traffic signs, lane markings), specialized models remain superior. However, for handling novelty and improving human-AI interaction, the ”cost” of using a more general VLSeg model is justifiable.
@@ V-A2 Standardizing Evaluation for VLSeg in ITS
    ITS-Specific Vocabulary: Evaluation should use a standardized set of ITS-specific prompts, including both common categories (e.g., “car,” “pedestrian”) and more complex, compositional queries (e.g., “red car turning left,” “pedestrian crossing between parked vehicles”).
@@
    Robustness Evaluation: Performance under corruptions (e.g., motion blur, adverse weather) should be systematically evaluated using standardized corruption sets like Cityscapes-C or BDD100K-C.This proposed protocol directly addresses gaps in existing benchmarks. For instance, while the Cityscapes leaderboard focuses on mIoU across 19 predefined classes, our protocol mandates evaluation on compositional queries (e.g., ’the car turning left’) and small object IoU, which are vital for language-guided ITS but are not primary metrics in today’s leaderboards. Furthermore, requiring systematic evaluation on corruption sets like Cityscapes-C elevates robustness from a research topic to a mandatory benchmark criterion.
@@ VI Advanced and Emerging Topics
3D VLSeg presents unique challenges. Point clouds are sparse, unordered, and unstructured, making them fundamentally different from the dense, grid-like structure of images. Early work in 3D segmentation focused on adapting CNN-like architectures to operate on voxels or directly on points [80, 81]. More recently, the focus has shifted to grounding language in these 3D spaces. Models like LidarCLIP [82] learn to align text descriptions with entire 3D point clouds. Building on this, 3D VLSeg models aim to segment specific objects in the point cloud based on a text prompt. This is often achieved by projecting 2D image features (from multiple camera views) onto the 3D point cloud, creating a text-aware 3D representation. The system can then respond to queries like “segment the point cloud of the truck in front of us” or “highlight the curb on the right”. This allows for a much more intuitive and powerful way to interact with and understand 3D sensor data, which is essential for tasks like 3D object detection, motion prediction, and path planning [83, 84].
@@ VI-B Video and Temporal Consistency
Several approaches are being explored. One common method is to use optical flow to propagate masks from one frame to the next. However, this can be error-prone, especially with fast-moving objects or camera motion. A more robust approach, exemplified by models like XMem [61], is to use a memory-based architecture. In this paradigm, the model maintains a “memory” of past frames and their segmentation masks. When processing a new frame, it uses attention mechanisms to query this memory, allowing it to re-identify and maintain a consistent segmentation of objects over long video sequences. Language can be used to initialize the tracking (e.g., “start tracking the blue car”) and to re-identify objects if the tracking is lost (“where is the blue car now?”). As video foundation models become more powerful, we expect to see more end-to-end V-VLSeg models that can reason about actions and events over time [85, 86].
@@ VI-C Federated and Collaborative Learning
Training powerful segmentation models requires vast amounts of diverse data, which raises significant privacy concerns, especially when the data is collected from personal vehicles. Federated Learning (FL) is a machine learning paradigm that addresses this issue. Instead of pooling raw data in a central server, the central model is sent to individual vehicles (or “clients”). Each client updates the model locally using its own private data, and only the model updates (gradients or weights) are sent back to the server to be aggregated. This allows a global model to learn from the collective data of the entire fleet without any raw driving data ever leaving the vehicle [87, 88]. Applying FL to VLSeg in ITS is an active research area, focusing on challenges like communication efficiency and handling the non-IID (non-independently and identically distributed) nature of data from different vehicles [89].
A related concept is Collaborative (or Collective) Perception. In this scenario, vehicles and infrastructure (e.g., smart traffic lights) communicate with each other, sharing high-level perception information, such as segmentation masks or object detections. For example, a vehicle whose view is occluded by a large truck could receive segmentation data from another vehicle at a better vantage point, allowing it to “see” the pedestrian that is about to cross the street. This creates a more robust and complete understanding of the driving scene than any single agent could achieve alone. Research in this area focuses on what information to share, how to fuse it effectively, and how to ensure the communication is secure and reliable [90, 91]. Language can act as a powerful and efficient communication medium in these systems, where one agent could send a compressed, semantic message like “pedestrian crossing from your right” to another.